OpenStudy (anonymous):

Divide and simplify. P^6/P^13

OpenStudy (anonymous):

P^-7

OpenStudy (saifoo.khan):

\[\Large \frac{P^a}{P^b} \to P^{a-b}\]

OpenStudy (anonymous):

\[6-13=-7\] \[\frac{p^6}{P^{13}}=p^{-7}=\frac{1}{P^7}\]

OpenStudy (anonymous):

since when does \(\large \to\) mean equal?

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

Thanks guys!